OpenStudy (anonymous):

In the triangle below, what is the length of the side opposite the 60 angle?

OpenStudy (anonymous):

\[A. 3\sqrt{3} \] \[B. \sqrt{3}\] \[C. 2\sqrt{3}\] \[D. 6\] 

OpenStudy (wolf1728):

The answer is D) 6 It is a 30-60-90 triangle and so the hypotenuse is twice the length of the shortest side.

OpenStudy (anonymous):

is not D